OpenStudy (anonymous):

What is the solution of n² – 49 = 0?

OpenStudy (campbell_st):

this is the difference of 2 squares \[n^2 - 49 = (n - 7)(n + 7) = 0\] now solve the binomials.

OpenStudy (anonymous):

n=7,-7

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

ok